OpenStudy (anonymous):

2 1/3m= 7/12? How would I solve this

OpenStudy (hockeychick23):

ok so the first thing you should do is make 2 1/3 into an improper fraction, do you know how to do that?

OpenStudy (anonymous):

Yes would it be 7/3?

OpenStudy (hockeychick23):

Correct, so then your equation will look like this: 7/3m=7/12. So the next thing you should do to get y by itself is multiply both side of the equation by the reciprocal of 7/3 to get m alone, The reciprocal of 7/3 is 3/7. 3/7(7/3m=7/12) so what is 3/7*7/12?

OpenStudy (anonymous):

21/84?

OpenStudy (hockeychick23):

Correct, and they actually have a common factor, so simplified it is 1/4. So your final answer is m= 1/4

OpenStudy (anonymous):

Thank you!

OpenStudy (hockeychick23):

No problem(: